OpenStudy (anonymous):

The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (anonymous):

Hint: surface area is two-dimensional.

OpenStudy (anonymous):

1) Surface Area = 2Lw + 2wh + 2Lh = 150cm² where: L = length w = width h = height So, if you new sides are trippled: Surface Area = 2(3L)(3w) + 2(3w)(3h) + 2(3L)(3h) = 18Lw + 18wh + 18Lh Set up the ratio of new surface area to old: (18Lw + 18wh + 18Lh)/(2Lw + 2wh + 2Lh) = Area/150 and solve for Area.

OpenStudy (ajanijones):

I think its 1,800

OpenStudy (anonymous):

i hope this helps.

OpenStudy (anonymous):

So I just add all that stuff up?